A bullet of mass 0.014 kg and initial speed 440 m/s penetrates an initially stationary wooden block with a of mass 0.44 kg and lodges in it.

How much mechanical energy was lost during the collision? (This mechanical energy was transformed into sound and friction generated heat.)
Answer #1

m1*u1 = (m1+m2)*V


V = (0.014*440)/(0.014+0.44) = 13.57 m/s


K1 = 0.5*m1*u1^2 = (0.5*0.014*440*440) = 1355.2 J

k2 = 0.5*(m1+m2)*V^2 = 0.5*(0.014+0.44)*13.57*13.57 = 41.8008923


loss = k1 - k2 = 1313.399 J
Answer #2

MEchanical energy lost = 1/2*M*V^2 = 1355.2 J
Answer #3

From conservation of momentum...

mv = (M+m)v

(.014)(440) = (.454)(v)

v = 13.6 m/s

Then the loss of energy is KE - KE

That is --> .5(.014)(440)2 - .5(.454)(13.6)2 = 1313.4 J lost